OpenStudy (anonymous):

Let x be the smaller one of two consecutive integers. If the sum of the squares of the two integers is less than three times the product of the two integers by 1, then A. x^(2)+(x+1)^(2)=3x(x+1)-1 B. x^(2)+(x+1)^(2)=3x(x+1)+1 C. 3[x^(2)+(x+1)^(2)]=x(x+1)-1 D. 3[x^(2)+(x+1)^(2)]=x(x+1)+1

hartnn (hartnn):

x is smaller so other = x+1

OpenStudy (anonymous):

I know

hartnn (hartnn):

where are u stuck ?

hartnn (hartnn):

sum of the squares of the two integers = x^(2)+(x+1)^(2)

OpenStudy (anonymous):

I know also

hartnn (hartnn):

three times the product of the two integers = 3x(x+1)

hartnn (hartnn):

sum of squares is less by 1 so x^(2)+(x+1)^(2)=3x(x+1)-1

OpenStudy (anonymous):

I am stuck in whether three times.... confusing.......wait a moment

hartnn (hartnn):

three times the product of the two integers = 3x(x+1)

OpenStudy (anonymous):

less than three times the product of the two integers by 1.... ok i got it.

OpenStudy (anonymous):

so I got the answer A correctly ;D

hartnn (hartnn):

good :)`

OpenStudy (jiteshmeghwal9):

If 'x' is smaller of two consecutive integers. Then, Larger integer=(x+1) A/q, => (x)^2+(x+1)^2=3(x)(x+1)-1

OpenStudy (anonymous):

Let x be the larger one of two consecutive odd numbers. If the sum of the squares of the two odd numbers is less than four times the product of the two odd numbers by 2, then x^(2)+(x-2)^(2)=4x(x-2)-2 isn't it?

hartnn (hartnn):

yes

OpenStudy (anonymous):

yeah XD thank you~

OpenStudy (jiteshmeghwal9):

you r correct @kryton1212 but i don't know that this is ur second question or first :/

hartnn (hartnn):

sorry, thats correct

OpenStudy (anonymous):

another question: Let k be a constant. Find the range of values of k such that the quadratic equation x^(2)+6x+k=3 has no real roots. @jiteshmeghwal9 there are two questions and are separated :) @hartnn never mind:)

hartnn (hartnn):

for no real roots D<0

hartnn (hartnn):

D = b^2-4ac = ?

OpenStudy (jiteshmeghwal9):

ohh ! okay @kryton1212 :)

OpenStudy (anonymous):

@hartnn D=36-4k

hartnn (hartnn):

36-4k <0

OpenStudy (anonymous):

k>9

hartnn (hartnn):

36-4k<0 -4k <-36 4k>36 yes k>9 is correct :)

OpenStudy (anonymous):

you are so powerful .... thank you so much ;D

hartnn (hartnn):

your welcome ^_^

OpenStudy (anonymous):

and the following questions are about surds....

OpenStudy (anonymous):

\[\frac{ 1 }{ 1+\sqrt{2} } +\frac{ 1 }{ \sqrt2+\sqrt{3} }+\frac{ 1 }{ \sqrt3+\sqrt{4} }+\frac{ 1 }{ \sqrt4+\sqrt{5} }=?\]

hartnn (hartnn):

wow..

hartnn (hartnn):

that will simplify greatly!

OpenStudy (jiteshmeghwal9):

``` MATH PROCESSING ERROR ``` ??

hartnn (hartnn):

no its ok

OpenStudy (jiteshmeghwal9):

But in my computer it's nt okay

hartnn (hartnn):

refresh the page

OpenStudy (anonymous):

I got the answer -1+sqrt5 but I don't know the way to do this...

hartnn (hartnn):

if u don't know then how did u get ? :P

OpenStudy (anonymous):

there are five choices in MC question. I calculate the question and answers:P

hartnn (hartnn):

multiply and divide by conjugates